Solve for the volume generated by revolving the given plane area about the given line using the circular ring method.

10. Within y = x^2 , y = 4x − x^2 ; about y = 6
Answer #1

1 -x disk about = 나가 기지 G-6) method din = 6-y,= 6-42-기는 dout = 6-y, = ? 2 지니가 키 A(5) - - not soul] -(1-td-t) (6)] 구름나기 계 것이 (